Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Monday, December 28, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 28, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/28/98 (Monday)

Segment 1 (Emidio to Copco Road): Report for 28 December 1998: Cutting of the 8" steel pipe along Copco Road resumed today. The sizing plate was run successfully through the short section near Cuddy Creek. Crews at that location then began preparing for the Hydrotest. Excavation, lowering in, and welding occurred at the Warbler Skip, as well as continuing dewatering into both a tank and an adjacent field. Excavation moved forward on Tejon Bench today. Crews continued to dig out the large bore exit at the base of Tejon Bench. Backfill crews moved forward with their work along Grapevine Road. The bore under Grapevine Creek continued. Fiber optic splicing was performed along Road B. Earthmoving began on the south side of the California Aqueduct crossing, as well as coating on the pipe. Crews on the north side of the aqueduct focused on establishing pressure on the pipe during the Hydrotest. A Non Compliance was issued for the violation of A-2, PC 1 and B-W-3, PC 1 for continued use of an unauthorized access road, outside of the ROW in San Joaquin kit fox habitat. A project memo was issued regarding the need for a qualified biologist to be present during earthmoving activities in kit fox habitat. Please note a correction for 12/27/98: Crews did not complete removing 8" pipe from Copco Road as stated.

Segment 4B (Lynwood to El Segundo): Report for 12-28-98. Station work continued at both El Segundo and Lynwood sites. Crews at Lynwood station progressed toward completing the block wall enclosing the station and with the final electrical tasks for the site. At the El Segundo station crews conducted a general clean up of the area and continued electrical work while a commission crews inspected the valves of the station.

Segment 5 (Lynwood to Wilmington): Report for 12-28-98. Work was delayed at the Edison site with plans for crews to begin work on the 29th of Dec. Near Tosco along Alameda St crews backfilled the bore pit running under the roadway, x-rayed a test head and removed a pig from the line. Meanwhile, at the Texaco site, crews worked to complete the tie-in under the roadway. Crews working at the Tosco station poured red concrete over electrical conduit, installed a meter valve and continued with general electrical work.

Inspection Summar

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page